OpenStudy (lolokay1):

I need help on a project, I understand some of it but I need some help please :)

OpenStudy (lolokay1):

1.A triangular section of a lawn will be converted to river rock instead of grass. Maurice insists that the only way to find a missing side length is to use the Law of Cosines. Johanna exclaims that only the Law of Sines will be useful. Describe a scenario where Maurice is correct, a scenario where Johanna is correct, and a scenario where both laws are able to be used. Use complete sentences and example measurements when necessary. 2.An archway will be constructed over a walkway. A piece of wood will need to be curved to match a parabola. Explain to Maurice how to find the equation of the parabola given the focal point and the directrix. 3.There are two fruit trees located at (3,0) and (−3, 0) in the backyard plan. Maurice wants to use these two fruit trees as the focal points for an elliptical flowerbed. Johanna wants to use these two fruit trees as the focal points for some hyperbolic flowerbeds. Create the location of two vertices on the y-axis. Show your work creating the equations for both the horizontal ellipse and the horizontal hyperbola. Include the graph of both equations and the focal points on the same coordinate plane. 4.A pipe needs to run from a water main, tangent to a circular fish pond. On a coordinate plane, construct the circular fishpond, the point to represent the location of the water main connection, and all other pieces needed to construct the tangent pipe. Submit your graph. You may do this by hand, using a compass and straight edge, or by using a graphing software program. 5.Two pillars have been delivered for the support of a shade structure in the backyard. They are both ten feet tall and the cross sections of each pillar have the same area. Explain how you know these pillars have the same volume without knowing whether the pillars are the same shape.

OpenStudy (lolokay1):

I know one and two but I am not sure on the rest

pooja195 (pooja195):

Hi there! Welcome To OpenStudy :-) I am unsure of these questions as I have not learned them yet however a better way to get help might be to post the questions one by one that way users don't feel intimitated by all the words :P I'll tag users who might know this stuff :) @mathmate @zepdrix

pooja195 (pooja195):

@jim_thompson5910

pooja195 (pooja195):

@skullpatrol :P

OpenStudy (lolokay1):

thank you so much! I just wanted them to have all the info on that problems :D

pooja195 (pooja195):

It's alright :) trying to see if people will help :P hopefully someone will come by

OpenStudy (lolokay1):

I hope so ! I have been stuck on this for a while :(

jimthompson5910 (jim_thompson5910):

`1.A triangular section of a lawn will be converted to river rock instead of grass. Maurice insists that the only way to find a missing side length is to use the Law of Cosines. Johanna exclaims that only the Law of Sines will be useful. Describe a scenario where Maurice is correct, a scenario where Johanna is correct, and a scenario where both laws are able to be used. Use complete sentences and example measurements when necessary.` Let's say we have a triangle ABC A = first angle B = second angle C = third angle opposite of angle A is side 'a' opposite of angle B is side 'b' opposite of angle C is side 'c' If we know two sides (say a and b) and the angle between them (angle C), then we must use the law of cosines to find side c. The law of cosines applies to SAS triangles If we know 2 angles, then we can easily find that missing angle. We would also need to know at least 1 side. With this collective info, we can use the law of sines. The law of sines is handy for AAS or ASA triangles. However, be careful that you may get trouble dealing with SSA triangles

jimthompson5910 (jim_thompson5910):

with the SSA case, 3 possibilities could occur * one triangle is possible * two triangles are possible * no triangles are possible

OpenStudy (lolokay1):

oh thank you so much but I got number one already haha but thanky uo

OpenStudy (lolokay1):

I need help with 3-5

jimthompson5910 (jim_thompson5910):

just for future reference, you should post each question as a separate thread. One problem per post. This is to avoid clutter and lag and such

jimthompson5910 (jim_thompson5910):

`3.There are two fruit trees located at (3,0) and (−3, 0) in the backyard plan. Maurice wants to use these two fruit trees as the focal points for an elliptical flowerbed. Johanna wants to use these two fruit trees as the focal points for some hyperbolic flowerbeds. Create the location of two vertices on the y-axis. Show your work creating the equations for both the horizontal ellipse and the horizontal hyperbola. Include the graph of both equations and the focal points on the same coordinate plane.` if the two foci are (3,0) and (-3,0) then what is the center? Do you know how to find the center?

OpenStudy (lolokay1):

again im so sorry I just wanted everyone to know the whole thing.

jimthompson5910 (jim_thompson5910):

I understand. It's just a lot to go through. Especially since each problem is pretty lengthy on its on.

OpenStudy (lolokay1):

b^2=a^2-c^2 ?

jimthompson5910 (jim_thompson5910):

hint: the center is the midpoint of the foci

OpenStudy (lolokay1):

wait no midpoint formula!

jimthompson5910 (jim_thompson5910):

do you know the midpoint formula?

OpenStudy (lolokay1):

not by heart but I can look it up

OpenStudy (lolokay1):

x1 + x2, y1+y2 both over 2

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

basically you add up the corresponding coordinates, then divide by 2

OpenStudy (lolokay1):

so the midpoint is 0

jimthompson5910 (jim_thompson5910):

you should get (0,0)

OpenStudy (lolokay1):

yes I did woohoo!

OpenStudy (lolokay1):

so I just have to plot those on a graph?

jimthompson5910 (jim_thompson5910):

yeah and an ellipse and hyperbola as well

jimthompson5910 (jim_thompson5910):

your teacher also wants the equations of the ellipse and hyperbola

OpenStudy (lolokay1):

do you mind helping me with those I am sorry

jimthompson5910 (jim_thompson5910):

where do you want the vertex point to be? pick any point on the x axis that is to the right of (3,0)

OpenStudy (lolokay1):

1,2?

jimthompson5910 (jim_thompson5910):

any point on the x axis is of the form (x,0)

OpenStudy (lolokay1):

ohh im dumb I sorry how about 2,0

jimthompson5910 (jim_thompson5910):

pick something to the right of (3,0)

OpenStudy (lolokay1):

oh sorry 4,0

jimthompson5910 (jim_thompson5910):

the two foci are (-3,0) and (3,0) the center is (0,0) the vertex you picked is (4,0)

jimthompson5910 (jim_thompson5910):

how far is it from the center to the vertex?

OpenStudy (lolokay1):

4,0

jimthompson5910 (jim_thompson5910):

a distance is a single number

OpenStudy (lolokay1):

oh so 4?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

ok so the ellipse is wider than it is tall, which means that the co-vertex will be some point on the y axis that is closer than 4 units to the center let's say the co-vertex is (0,2)

jimthompson5910 (jim_thompson5910):

the two foci are (-3,0) and (3,0) the center is (0,0) vertices (4,0) and (-4,0) co-vertices (0,2) and (0,-2)

jimthompson5910 (jim_thompson5910):

are you able to take that info and form an equation for the ellipse?

OpenStudy (lolokay1):

yes!

jimthompson5910 (jim_thompson5910):

ok great

OpenStudy (lolokay1):

x^ over a^ + y^2 over b^2 =1 correct?

jimthompson5910 (jim_thompson5910):

Did you mean this? \[\Large \frac{(x-h)^2}{a^2} + \frac{(y-k)^2}{b^2} = 1\]

jimthompson5910 (jim_thompson5910):

General equation of an ellipse \[\Large \frac{(x-h)^2}{a^2} + \frac{(y-k)^2}{b^2} = 1\] (h,k) is the center 'a' determines how wide the ellipse is 'b' determine how tall the ellipse is

OpenStudy (lolokay1):

yes:)

jimthompson5910 (jim_thompson5910):

In this case, the center is (0,0) so the 'h' and 'k' will effectively go away

jimthompson5910 (jim_thompson5910):

making \[\Large \frac{(x-h)^2}{a^2} + \frac{(y-k)^2}{b^2} = 1\] turn into \[\Large \frac{x^2}{a^2} + \frac{y^2}{b^2} = 1\] now you just need to figure out what 'a' and 'b' are

OpenStudy (lolokay1):

So what do I do D:

jimthompson5910 (jim_thompson5910):

In this case, the distance from the center (0,0) and the vertex (4,0) is 4 units so a = 4

jimthompson5910 (jim_thompson5910):

the distance from center to focus is 3 units so c = 3

OpenStudy (lolokay1):

ohhhhhh okay

OpenStudy (lolokay1):

so now b

jimthompson5910 (jim_thompson5910):

http://www.mathwords.com/f/foci_ellipse.htm a^2 = b^2 + c^2 b^2 = a^2 - c^2 b^2 = 4^2 - 3^2 b^2 = ???

jimthompson5910 (jim_thompson5910):

btw we don't need to find the value of b, we can stop at b^2

OpenStudy (lolokay1):

sorry my computers lagging so I cant open the link

OpenStudy (lolokay1):

so I just put that as the answer to the ellipse?

jimthompson5910 (jim_thompson5910):

you should find that b^2 = 7

jimthompson5910 (jim_thompson5910):

a = 4 ---> a^2 = 16 b^2 = 7 so \[\Large \frac{x^2}{a^2} + \frac{y^2}{b^2} = 1\] becomes \[\Large \frac{x^2}{16} + \frac{y^2}{7} = 1\] which is the equation of the ellipse with these properties * center at (0,0) * foci at (-3,0) and (3,0) * vertices at (4,0) and (-4,0)

OpenStudy (lolokay1):

I am so sorry I am so confused

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (lolokay1):

from right after I said so now b

jimthompson5910 (jim_thompson5910):

did you see the attachment where I have a screenshot of the formulas needed? in this case, we use a^2 = b^2 + c^2

OpenStudy (lolokay1):

ohhh is that the ellipse formula?

jimthompson5910 (jim_thompson5910):

it will help figure out the ellipse equation. It's not the actual ellipse equation

jimthompson5910 (jim_thompson5910):

I'll be right back

OpenStudy (lolokay1):

ohhhh I see thank you! I just didn't know how you got that